5. What quantity of heat is needed to melt 120 g of ice at -20 o

c to water at 30 o

c? Specific 

heat capacity of water and ice are 4200 JKg-1K

-1 and 2100 JKg-1K

-1

respectively. And 

also to specific latent heat 0f ice as 3.36 x 105

JK-1


1
Expert's answer
2020-07-08T13:15:12-0400
"Q=m(L+c_i(T_1-T)+c_w(T_2-T_1))\\Q=0.12(3.36\cdot10^5+2100(20)+4200(30))\\Q=60.5\ kJ"